1. Andrew has a constant elasticity of substitution (CES) utility function,

U(x1, x2) =  where  and


Determine Andrew’s optimal bundle (x1, x2) in terms of his income m and prices of the two goods, p1 and p2.

 

 

2. Lynn has a Cobb-Douglas utility function 

U(x1, x2) =


What share of her budget does she spend on x1 (recorded music tracks) and x2 (live music) in terms of her income m = $30, prices of the two goods, p1 = $0.5 and p2= $1?


3. Celine’s quasilinear utility function is



Her budget for these two goods is $10. Originally the prices are p1 = p2 = $1. However, the price of the first good rises to $2. Determine the substitution, income and total effect of this price change on the demand for x1.


1
Expert's answer
2021-10-06T09:42:58-0400

1.

Given in the question-

Income = M

Price of good 1 = P1

Price of good 2 = P2

Utility function "U(x_1, x_2) = x_1 x_2"

General Budget equation =

"P_1x_1 + P_2x_2 = M ................(1)"

At equilibrium -

"MRS = \frac{P_1}{p_2}\\

MRS =\frac{ x_2}{x_1}"


"\frac{x_2}{x_1 }=\frac{ P_1}{p2}\\

x_2 = \frac{p_1x_1}{p_2}\\

Putting \space value\space in\space equation \space "1"\\

p_1x_1 + p_2x_2 =M\\

p_1x_1 + p_2\ times (\ frac{p_1x_1}{p_2}) = M

p_1x_1 + p_1x_1 = M\\

2p_1x_1 = M\\

x_1 = \frac{M}{2P_1}\\

similarly,\\

x_2 = \frac{M}{2P_2}\\

optimal \space bundles \space will\space be \space "\\frac{M}{2P_1 }, \\frac{M}{2P_2}""

2.

Given in the question-

Income = M

Price of good 1 = P1 =0.5

Price of good 2 = P2 =1

Utility function "U(x_1, x_2) = x_1 x_2"

General Budget equation =

"P_1x_1 + P_2x_2 = M ................(1)\\30=0.5X_1+X_2"

At equilibrium -

"MRS = \frac{P_1}{p_2}\\MRD=\frac{0.5}{1}
=0.5\\
MRS =\frac{ x_2}{x_1}\frac{1}{0.5}=2"

3.

"\frac{x_2}{x_1 }=\frac{ P_1}{p2}\\

x_2 = \frac{p_1x_1}{p_2}\\

Putting \space value\space in\space equation \space "1"\\

p_1x_1 + p_2x_2 =M\\

p_1x_1 + p_2\ times (\ frac{p_1x_1}{p_2}) = M

p_1x_1 + p_1x_1 = M\\

2p_1x_1 = M\\

x_1 = \frac{M}{2P_1}\\

similarly,\\

x_2 = \frac{M}{2P_2}\\

optimal \space bundles \space will\space be \space "\\frac{M}{2P_1 }, \\frac{M}{2P_2}""

Therefore